For the function f(x), f(2.3)= -7.2 and f(8.4)= 9.1. What is the equation for the function f(x)?
Answer by stanbon(75887)

For the function f(x), f(2.3)= -7.2 and f(8.4)= 9.1. What is the equation for the function f(x)?
-----
slope = (9.1--7.2)/(8.4-2.3) = 16.3/6.1 = 163/61
--------
Form: y = mx + b
Solve for "b":
9.1 = (163/61)8.4 + b
b = 9.1 - 22.45 = -13.35
------
Equation:
y = (163/61)x - 13.35
=======================
Cheers,
Stan H.
================